OpenStudy (anonymous):

solve each equation. show work please.its a fraction just dnt know how to type on here 7n+(-1)/8 =-42

OpenStudy (anonymous):

7n-(1/8)=-42...multiply the entire equation by 8 to get rid of the denominator...you get...56n-1=-336...add 1 to 336...56n=-335...Finally n=-335/56 or around 5.98

jhonyy9 (jhonyy9):

56n-1=-336 56n=-335 n=-335/56